# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 4.2.2 CV-STEM Formulation for State Estimation
If (4.13) of Theorem 4.3 holds, then we have the following bound:
@@
Finally, the CV-STEM estimation framework is summarized in Theorem 4.5 Tsukamoto and Chung [2021, 2021], Tsukamoto et al. [2021], Tsukamoto and Chung [2019].
@@
Suppose that $\alpha$, $\alpha_{G}$, $\bar{d}_{e0}$, $\bar{d}_{e1}$, $\bar{g}_{e0}$, $\bar{g}_{e1}$, and $L_{m}$ in (4.13) and (4.19) are given. If the pair $(A,C)$ is uniformly observable, the non-convex optimization problem of minimizing the upper bound (4.19) is equivalent to the following convex optimization problem with the contraction constraint (4.13) and $\mathrm{I}\preceq\bar{W}\preceq\chi\mathrm{I}$ of (3.26):
@@
where $c_{0}$, $c_{1}$, and $s$ are as defined in (4.19) of Theorem 4.4, $c_{2}\in\mathbb{R}_{\geq 0}$, and $P$ is some performance-based cost function as in Theorem 4.2.
The weight $c_{1}$ for $\nu^{s}$ indicates how much we trust the measurement $y(t)$. Using non-zero $c_{2}$ enables finding contraction metrics optimal in a different sense in terms of $P$. Furthermore, the coefficients of the SDC parameterizations $\varrho_{a}$ and $\varrho_{c}$ in Lemma 3.1, i.e., $A=\sum\varrho_{a,i}A_{i}$ and $C=\sum\varrho_{c,i}C_{i}$ in (4.5) and (4.6), can also be treated as decision variables by convex relaxation Dani et al. [2015], thereby adding a design flexibility to mitigate the effects of external disturbances while verifying the system observability.
@@
The proposed optimization problem is convex as its objective and constraints are convex in terms of decision variables $\chi$, $\nu$, and $\bar{W}$ (see Remark 4.1). Also, larger $\bar{d}_{e1}$ and $\bar{g}_{e1}$ in (4.3) and (4.4) imply larger measurement uncertainty. Thus by definition of $c_{1}$ in Theorem 4.4, the larger the weight of $\nu$, the less confident we are in $y(t)$ (see Example 4.2). The last statement on the SDC coefficients for guaranteeing observability follows from Proposition 1 of Dani et al. [2015] and Proposition 1 of Tsukamoto and Chung [2021]. $\qed$
@@
The weights $c_{0}$ and $c_{1}$ of the CV-STEM estimation of Theorem 4.5 has an analogous trade-off to the case of the Kalman filter with the process and sensor noise covariance matrices, $Q$ and $P$, respectively, since the term $c_{0}\chi$ in upper bound of the steady-state tracking error in (4.19) becomes dominant if measurement noise is much smaller than process noise ($\bar{d}_{e0}\gg\bar{d}_{e1}$ or $\bar{g}_{e0}\gg\bar{g}_{e1}$), and the term $c_{1}\nu^{s}$ becomes dominant if measurement noise is much greater than process noise ($\bar{d}_{e0}\ll\bar{d}_{e1}$ or $\bar{g}_{e0}\ll\bar{g}_{e1}$). In particular Tsukamoto and Chung [2021], Tsukamoto et al. [2021],
@@
This is also because the solution $Q=P^{-1}\succ 0$ of the Kalman filter Riccati equation [Brown and Hwang, 1997, p. 375], $\dot{P}=AP+PA^{\top}-PC^{\top}R^{-1}CP+Q$, can be viewed as a positive definite matrix that defines a contraction metric as discussed in Example 2.3.
4.3 Control Contraction Metrics (CCMs)
As briefly discussed in Remark 3.4, the concept of a CCM Manchester and Slotine [2017], Singh et al. [2017, 2020], Wang et al. [2019, 2020], Manchester and Slotine [2018] is introduced to extend contraction theory to design differential feedback control laws for control-affine deterministic nonlinear systems (3.1), $\dot{x}=f(x,t)+B(x,t)u$. Let us show that the CCM formulation could also be considered in the CV-STEM control of Theorem 4.2, where similar ideas have been investigated in Tsukamoto et al. [2021], Singh et al. [2017], Zhao et al. [2021].
@@ 4.3 Control Contraction Metrics (CCMs)
Suppose the CV-STEM control of Theorem 4.2 is designed with its contraction condition replaced by the following set of convex constraints along with $\mathrm{I}\preceq\bar{W}\preceq\chi\mathrm{I}$ of (3.26):
@@
Given a CCM, there exists a differential feedback controller $\delta u=k(x,\delta x,u,t)$ that stabilizes the following differential dynamics of (3.1) with $d_{c}=0$ along all solutions (i.e., the closed-loop dynamics is contracting as in Definition 2.3):
@@
where $A=\partial f/\partial x+\sum_{i=1}^{m}(\partial b_{i}/\partial x)u_{i}$ and $\delta u$ is a tangent vector to a smooth path of controls at $u$. Furthermore, it can be computed as follows Manchester and Slotine [2017], Singh et al. [2017], Manchester and Slotine [2018]:
@@
where $\partial_{\mu}\gamma=\partial\gamma/\partial\mu$, $\gamma$ is the minimizing geodesic with $\gamma(0,t)=x_{d}$ and $\gamma(1,t)=x$ for $\mu\in[0,1]$, $(x_{d},u_{d})$ is a given target trajectory in $\dot{x}_{d}=f(x_{d},t)+B(x_{d},t)u_{d}$ of (3.3), and the computation of $k$ is given in Manchester and Slotine [2017], Singh et al. [2017] (see Remark 4.2). Furthermore, Theorem 2.4 for deterministic disturbance rejection still holds and the CCM controller (4.25) thus possesses the same sense of oplimality as for the CV-STEM control in Theorem 4.2.
@@
where $A=\partial f/\partial x+\sum_{i=1}^{m}(\partial b_{i}/\partial x)u_{i}$ for the differential dynamics (4.24). The relation (4.26) states that $\delta x$ orthogonal to the span of actuated directions $b_{i}$ is naturally contracting, thereby implying the stabilizability of the system (3.1) with $d_{c}=0$, i.e., $\dot{x}=f(x,t)+B(x,t)u$. See Manchester and Slotine [2017], Singh et al. [2017], Manchester and Slotine [2018] (and Tsukamoto et al. [2021] on the CV-STEM formulation) for the rest of the proof. $\qed$
@@
Let us consider a case where $k$ of the differential feedback controller does not depend on $u$, and is defined explicitly by $M(x,t)\succ 0$ as follows Tsukamoto et al. [2021]:
@@
yielding the convex optimization-based control synthesis algorithm of Theorem 4.6 independently of $(x_{d},u_{d})$ similar to that of Theorem 4.2 (see Tsukamoto et al. [2021] for details).
@@
If $B$ is of the form $0,\mathrm{I}]^{\top}$ for the zero matrix $0\in\mathbb{R}^{n_{1}\times m}$ and identity matrix $\mathrm{I}\in\mathbb{R}^{n_{2}\times m}$ with $n=n_{1}+n_{2}$, the condition (4.29) says that $M$ should not depend on the last $n_{2}$ state variables Manchester and Slotine [2017].
@@
We could consider stochastic perturbation in Theorem 4.6 using Theorem 2.5, even with the differential control law of the form (4.25) or (4.27) as demonstrated in Tsukamoto et al. [2021]. Also, although the relation (4.23) or (4.29) is not included as a constraint in Theorem 4.2 for simplicity of discussion, the dependence of $\dot{\bar{W}}$ on $u$ in Theorem 4.2 can be removed by using it in a similar way to Tsukamoto et al. [2021].
@@
As stated in (4.25), the computation of the differential feedback gain $k(x,\delta x,u,t)$ and minimizing geodesics $\gamma$ is elaborated in Manchester and Slotine [2017], Singh et al. [2017]. For example, if $M$ is state-independent, then geodesics are just straight lines.
@@
Let us again emphasize that, as delineated in Sec. 3.1, the differences between the SDC- and CCM-based CV-STEM frameworks in Theorems 4.2 and 4.6 arise only from their different form of controllers in $u=u_{d}-K(x-x_{d})$ of (3.15) and $u=u_{d}+\int^{1}_{0}kd\mu$ of (4.25), leading to the trade-offs outlined in Table 3.
4.4 Remarks in CV-STEM Implementation
@@
We propose some useful techniques for the practical application of the CV-STEM in Theorems 4.24.5, and 4.6. Note that the CV-STEM requires solving a convex optimization problem at each time instant, but its solution can be approximated with formal stability guarantees to enable faster computation using machine learning techniques as shall be seen in Sec. 6.
4.4.1 Performance-based Cost Function
@@
Selecting $c_{2}=0$ in Theorems 4.2, 4.5, and 4.6 yields a convex objective function that allows for a systematic interpretation of its weights as seen earlier in Examples 4.1 and 4.2. One could also select $c_{2}>0$ to augment the CV-STEM with other control and estimation performances of interest, as long as $P(\nu,\chi,\bar{W})$ in (4.2) and (4.21) is convex. For example, we could consider a steady-state tracking error bound of parametric uncertain systems as $P$, using the adaptive control technique to be discussed in Sec. 8 Tsukamoto et al. [2021]. Such a modification results in a contraction metric optimal in a different sense.
4.4.2 Selecting and Computing CV-STEM Parameters
@@
The CV-STEM optimization problems derived in Theorems 4.24.5, and 4.6 are convex if we assume that $\alpha$, $\alpha_{G}$, and $L_{m}$ are given. However, these parameters would also affect the optimality of resultant contraction metrics. In Tsukamoto and Chung [2021], Tsukamoto et al. [2021], Singh et al. [2017], a line search algorithm is performed to find optimal $\alpha$ and $\alpha_{G}$, while the Lipschitz constraint given with $L_{m}$ is guaranteed by spectrally-normalization Miyato et al. [2018], Tsukamoto et al. [2021] as shall be seen in detail in Sec. 6.3 (see Revay et al. [2020, 2021], Manchester et al. [2021] for contraction theory-based techniques for obtaining Lipschitz bounds). Also, the CV-STEM can be formulated as a finite-dimensional problem by using backward difference approximation on $\dot{\bar{W}}$, where we can then use $\bar{W}\preceq-\mathrm{I}$ to get a sufficient condition of its constraints, or we could alternatively solve it along pre-computed trajectories $\{x(t_{i})\}_{i=0}^{M}$ as in Tsukamoto and Chung [2021]. In Sec. 6, we use a parameterized function such as neural networks Hornik et al. [1989], Cybenko [1989], Funahashi [1989] for approximating $M$ to explicitly compute $\dot{M}$.
@@
5 Contraction Theory for Learning-based Control
Machine learning techniques, e.g., reinforcement learning Sutton and Barto [1998], Bertsekas and Tsitsiklis [1996], Everett et al. [2018], Berkenkamp et al. [2017], imitation learning Carius et al. [2020], Rivière et al. [2020], Ho and Ermon [2016], Gupta et al. [2018], Kuefler et al. [2017], and neural networks Hornik et al. [1989], Cybenko [1989], Funahashi [1989], have gained popularity due to their ability to achieve a large variety of innovative engineering and scientific tasks which have been impossible heretofore. Starting from this section, we will see how contraction theory enhances learning-based and data-driven automatic control frameworks providing them with formal optimality, stability, and robustness guarantees. In particular, we present Theorems 5.2 and 5.3 for obtaining robust exponential bounds on trajectory tracking errors of nonlinear systems in the presence of learning errors, whose steady-state terms are again written as a function of the condition number $\chi=\overline{m}/\underline{m}$ of a positive definite matrix $M$ that defines a contraction metric, consistently with the CV-STEM frameworks of Sec. 4.
5.1 Problem Formulation
Let us consider the following virtual nonlinear system as in Theorem 2.2:

### v8

@@ 4.2.2. CV-STEM Formulation for State Estimation
Convexity of Steady-State Estimation Error BoundsThm:CV-STEMobjective-est If (4.13) of Theorem LABEL:THM:Thm:CVSTEM-LMI-est holds, then we have the following bound:
@@
Finally, the CV-STEM estimation framework is summarized in Theorem LABEL:THM:Thm:CV-STEM-estimation mypaperTAC, ncm, nscm, mypaper.
@@
Robust Nonlinear Estimation via Convex OptimizationThm:CV-STEM-estimation Suppose that $\alpha$, $\alpha_{G}$, $\bar{d}_{e0}$, $\bar{d}_{e1}$, $\bar{g}_{e0}$, $\bar{g}_{e1}$, and $L_{m}$ in (4.13) and (4.19) are given. If the pair $(A,C)$ is uniformly observable, the non-convex optimization problem of minimizing the upper bound (4.19) is equivalent to the following convex optimization problem with the contraction constraint (4.13) and $\mathrm{I}\preceq\bar{W}\preceq\chi\mathrm{I}$ of (3.26):
@@
where $c_{0}$, $c_{1}$, and $s$ are as defined in (4.19) of Theorem LABEL:THM:Thm:CV-STEMobjective-est, $c_{2}\in\mathbb{R}_{\geq 0}$, and $P$ is some performance-based cost function as in Theorem LABEL:THM:Thm:CV-STEM.
The weight $c_{1}$ for $\nu^{s}$ indicates how much we trust the measurement $y(t)$. Using non-zero $c_{2}$ enables finding contraction metrics optimal in a different sense in terms of $P$. Furthermore, the coefficients of the SDC parameterizations $\varrho_{a}$ and $\varrho_{c}$ in Lemma 3.1, i.e., $A=\sum\varrho_{a,i}A_{i}$ and $C=\sum\varrho_{c,i}C_{i}$ in (4.5) and (4.6), can also be treated as decision variables by convex relaxation Ref:Stochastic, thereby adding a design flexibility to mitigate the effects of external disturbances while verifying the system observability.
@@
The proposed optimization problem is convex as its objective and constraints are convex in terms of decision variables $\chi$, $\nu$, and $\bar{W}$ (see Remark 4.1). Also, larger $\bar{d}_{e1}$ and $\bar{g}_{e1}$ in (4.3) and (4.4) imply larger measurement uncertainty. Thus by definition of $c_{1}$ in Theorem LABEL:THM:Thm:CV-STEMobjective-est, the larger the weight of $\nu$, the less confident we are in $y(t)$ (see Example LABEL:EX:ex:kalmanexample). The last statement on the SDC coefficients for guaranteeing observability follows from Proposition 1 of Ref:Stochastic and Proposition 1 of mypaperTAC. $\qed$
@@
Kalman Filter Analogyex:kalmanexample The weights $c_{0}$ and $c_{1}$ of the CV-STEM estimation of Theorem LABEL:THM:Thm:CV-STEM-estimation have an analogous trade-off to the case of the Kalman filter with the process and sensor noise covariance matrices, $Q$ and $P$, respectively, since the term $c_{0}\chi$ in upper bound of the steady-state tracking error in (4.19) becomes dominant if measurement noise is much smaller than process noise ($\bar{d}_{e0}\gg\bar{d}_{e1}$ or $\bar{g}_{e0}\gg\bar{g}_{e1}$), and the term $c_{1}\nu^{s}$ becomes dominant if measurement noise is much greater than process noise ($\bar{d}_{e0}\ll\bar{d}_{e1}$ or $\bar{g}_{e0}\ll\bar{g}_{e1}$). In particular ncm, nscm,
@@
This is also because the solution $Q=P^{-1}\succ 0$ of the Kalman filter Riccati equation [Brown:680442, p. 375], $\dot{P}=AP+PA^{\top}-PC^{\top}R^{-1}CP+Q$, can be viewed as a positive definite matrix that defines a contraction metric as discussed in Example LABEL:EX:ex:lyapunov_contraction.
4.3. Control Contraction Metrics (CCMs)
As briefly discussed in Remark 3.4, the concept of a CCM ccm, 7989693, 47710, WANG201944, vccm, rccm is introduced to extend contraction theory to design differential feedback control laws for control-affine deterministic nonlinear systems (3.1), $\dot{x}=f(x,t)+B(x,t)u$. Let us show that the CCM formulation could also be considered in the CV-STEM control of Theorem LABEL:THM:Thm:CV-STEM, where similar ideas have been investigated in cdc_ncm, 7989693, zhao2021tubecertified.
@@ 4.3. Control Contraction Metrics (CCMs)
Control Contraction MetricsThm:ccm_cvstem Suppose the CV-STEM control of Theorem LABEL:THM:Thm:CV-STEM is designed with its contraction condition replaced by the following set of convex constraints along with $\mathrm{I}\preceq\bar{W}\preceq\chi\mathrm{I}$ of (3.26):
@@
Given a CCM, there exists a differential feedback controller $\delta u=k(x,\delta x,u,t)$ that stabilizes the following differential dynamics of (3.1) with $d_{c}=0$ along all solutions (i.e., the closed-loop dynamics is contracting as in Definition LABEL:DEF:Def:contraction):
@@
where $A=\partial f/\partial x+\sum_{i=1}^{m}(\partial b_{i}/\partial x)u_{i}$ and $\delta u$ is a tangent vector to a smooth path of controls at $u$. Furthermore, it can be computed as follows ccm, 7989693, rccm:
@@
where $\partial_{\mu}\gamma=\partial\gamma/\partial\mu$, $\gamma$ is the minimizing geodesic with $\gamma(0,t)=x_{d}$ and $\gamma(1,t)=x$ for $\mu\in[0,1]$, $(x_{d},u_{d})$ is a given target trajectory in $\dot{x}_{d}=f(x_{d},t)+B(x_{d},t)u_{d}$ of (3.3), and the computation of $k$ is given in ccm, 7989693 (see Remark 4.2). Furthermore, Theorem LABEL:THM:Thm:Robust_contraction_original for deterministic disturbance rejection still holds and the CCM controller (4.25) thus possesses the same sense of oplimality as for the CV-STEM control in Theorem LABEL:THM:Thm:CV-STEM.
@@
where $A=\partial f/\partial x+\sum_{i=1}^{m}(\partial b_{i}/\partial x)u_{i}$ for the differential dynamics (4.24). The relation (4.26) states that $\delta x$ orthogonal to the span of actuated directions $b_{i}$ is naturally contracting, thereby implying the stabilizability of the system (3.1) with $d_{c}=0$, i.e., $\dot{x}=f(x,t)+B(x,t)u$. See ccm, 7989693, rccm (and cdc_ncm on the CV-STEM formulation) for the rest of the proof. $\qed$
@@
Control Contraction Metrics I Let us consider a case where $k$ of the differential feedback controller does not depend on $u$, and is defined explicitly by $M(x,t)\succ 0$ as follows cdc_ncm:
@@
yielding the convex optimization-based control synthesis algorithm of Theorem LABEL:THM:Thm:ccm_cvstem independently of $(x_{d},u_{d})$ similar to that of Theorem LABEL:THM:Thm:CV-STEM (see cdc_ncm for details).
@@
Control Contraction Metrics II If $B$ is of the form $0,\mathrm{I}]^{\top}$ for the zero matrix $0\in\mathbb{R}^{n_{1}\times m}$ and identity matrix $\mathrm{I}\in\mathbb{R}^{n_{2}\times m}$ with $n=n_{1}+n_{2}$, the condition (4.29) says that $M$ should not depend on the last $n_{2}$ state variables ccm.
@@
We could consider stochastic perturbation in Theorem LABEL:THM:Thm:ccm_cvstem using Theorem LABEL:THM:Thm:robuststochastic, even with the differential control law of the form (4.25) or (4.27) as demonstrated in cdc_ncm. Also, although the relation (4.23) or (4.29) is not included as a constraint in Theorem LABEL:THM:Thm:CV-STEM for simplicity of discussion, the dependence of $\dot{\bar{W}}$ on $u$ in Theorem LABEL:THM:Thm:CV-STEM can be removed by using it in a similar way to cdc_ncm.
@@
As stated in (4.25), the computation of the differential feedback gain $k(x,\delta x,u,t)$ and minimizing geodesics $\gamma$ is elaborated in ccm, 7989693. For example, if $M$ is state-independent, then geodesics are just straight lines.
@@
Let us again emphasize that, as delineated in Sec. 3.1, the differences between the SDC- and CCM-based CV-STEM frameworks in Theorems LABEL:THM:Thm:CV-STEM and LABEL:THM:Thm:ccm_cvstem arise only from their different form of controllers in $u=u_{d}-K(x-x_{d})$ of (3.15) and $u=u_{d}+\int^{1}_{0}kd\mu$ of (4.25), leading to the trade-offs outlined in Table 3.
4.4. Remarks in CV-STEM Implementation
@@
We propose some useful techniques for the practical application of the CV-STEM in Theorems $LABEL:THM:Thm:CV-STEM$-estimation, and LABEL:THM:Thm:ccm_cvstem. Note that the CV-STEM requires solving a convex optimization problem at each time instant, but its solution can be approximated with formal stability guarantees to enable faster computation using machine learning techniques as shall be seen in Sec. 6.
4.4.1. Performance-based Cost Function
@@
Selecting $c_{2}=0$ in Theorems LABEL:THM:Thm:CV-STEM, LABEL:THM:Thm:CV-STEM-estimation, and LABEL:THM:Thm:ccm_cvstem yields a convex objective function that allows for a systematic interpretation of its weights as seen earlier in Examples LABEL:EX:ex:lqrexample and LABEL:EX:ex:kalmanexample. One could also select $c_{2}>0$ to augment the CV-STEM with other control and estimation performances of interest, as long as $P(\nu,\chi,\bar{W})$ in (4.2) and (4.21) is convex. For example, we could consider a steady-state tracking error bound of parametric uncertain systems as $P$, using the adaptive control technique to be discussed in Sec. 8 ancm. Such a modification results in a contraction metric optimal in a different sense.
4.4.2. Selecting and Computing CV-STEM Parameters
@@
The CV-STEM optimization problems derived in Theorems $LABEL:THM:Thm:CV-STEM$-estimation, and LABEL:THM:Thm:ccm_cvstem are convex if we assume that $\alpha$, $\alpha_{G}$, and $L_{m}$ are given. However, these parameters would also affect the optimality of resultant contraction metrics. In ncm, nscm, 7989693, a line search algorithm is performed to find optimal $\alpha$ and $\alpha_{G}$, while the Lipschitz constraint given with $L_{m}$ is guaranteed by spectrally-normalization miyato2018spectral, nscm as shall be seen in detail in Sec. 6.3 (see revay2020lipschitz, revay2021recurrent, cdc_systemid for contraction theory-based techniques for obtaining Lipschitz bounds). Also, the CV-STEM can be formulated as a finite-dimensional problem by using backward difference approximation on $\dot{\bar{W}}$, where we can then use $\bar{W}\preceq-\mathrm{I}$ to get a sufficient condition of its constraints, or we could alternatively solve it along pre-computed trajectories $\{x(t_{i})\}_{i=0}^{M}$ as in ncm. In Sec. 6, we use a parameterized function such as neural networks neural1, neural2, neural3 for approximating $M$ to explicitly compute $\dot{M}$.
@@
5. Contraction Theory for Learning-based Control
Machine learning techniques, e.g., reinforcement learning sutton, ndp, 8593871, NIPS2017_766ebcd5, imitation learning 9001182, glas, NIPS2016_cc7e2b87, 8578338, 7995721, and neural networks neural1, neural2, neural3, have gained popularity due to their ability to achieve a large variety of innovative engineering and scientific tasks which have been impossible heretofore. Starting from this section, we will see how contraction theory enhances learning-based and data-driven automatic control frameworks providing them with formal optimality, stability, and robustness guarantees. In particular, we present Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto for obtaining robust exponential bounds on trajectory tracking errors of nonlinear systems in the presence of learning errors, whose steady-state terms are again written as a function of the condition number $\chi=\overline{m}/\underline{m}$ of a positive definite matrix $M$ that defines a contraction metric, consistently with the CV-STEM frameworks of Sec. 4.
5.1. Problem Formulation
Let us consider the following virtual nonlinear system as in Theorem LABEL:THM:Thm:partial_contraction: